What is the probability that a boy will score lower than a girl?
There is insufficient information for us to answer this question. Please edit the question to include more context or relevant information. It would be useful to know what they were trying to score at!
What does a prognosis of fair mean?
While subjective, "fair" indicates a moderate probability of returning to normal function - it's not great odds, but the risk of severe complications or death is low.
Why probability always lays 0 to 1?
0 represents that is will never happen. It means that there is no probability that it will happen. 1 represents that is is certain to happen. You cannot be more certain than definitely certain, and something cannot be less probable than not at all.
How do you determine experimental probability?
You conduct the appropriate experiment repeatedly and calculate the proportion of times that the specified event occurs compared with the total number of experimental trials.
You conduct the appropriate experiment repeatedly and calculate the proportion of times that the specified event occurs compared with the total number of experimental trials.
You conduct the appropriate experiment repeatedly and calculate the proportion of times that the specified event occurs compared with the total number of experimental trials.
You conduct the appropriate experiment repeatedly and calculate the proportion of times that the specified event occurs compared with the total number of experimental trials.
What is the probability that the fourth toss is tails?
1/2. What the previous tosses are has no impact on what a fourth toss of a coin is, so the chance of getting a tails is the same on any throw.
75%
What are the odds of life occurring in the universe?
The probability is 1, since it has already happened. There has been life on earth for millions of years - long before humans.
How many permutations are in the word permutation?
239,500,800
12!/2!
* * * * *
Actually, as the word "permutation" [not permutations] has 11 letters, the answer is 11!/2! = 19,958,400
Why do sociologists sample a population they are going to study?
Because it would usually be prohibitively expensive and time consuming to collect information from every single member of the population and because the law of large numbers is sufficiently robust to allow most characteristics of a population to be measured with sufficient accuracy from a sample.
What happens if your app never loads?
it means you should re download it or your phone does not support the app
When you toss a coin the probability that it will land heads up is 25 percent?
No, when you toss a coin there is a 50 percent chance it will land heads up.
The probability of getting an even sum on two dice is 18 in 36 or 1 in 2 or 0.5. The probability of doing that three times in a row is 0.53 or 0.125.
Around 1%
What happens after arthroplasty?
Patients who have undergone arthroplasty must be careful not to over stress or destabilize the joint
How do you find p(b) when p(a) is 23 p(ba) is 12 and p(a U b) is 45 and is a dependent event?
There are symbols missing from your question which I cam struggling to guess and re-insert.
p(a) = 2/3
p(b ??? a) = 1/2
p(a ∪ b) = 4/5
p(b) = ?
Why use the set notation of Union on the third given probability whereas the second probability has something missing but the "sets" are in the other order, and the order wouldn't matter in sets.
There are two possibilities:
1) The second probability is: p(b ∩ a) = p(a ∩ b) = 1/2
→ p(a) + p(b) = p(a ∪ b) + p(a ∩ b)
→ p(b) = p(a ∪ b) + p(a ∩ b) - p(a)
= 4/5 + 1/2 - 2/3
= 24/30 + 15/30 - 20/30
= 19/30
2) The second and third probabilities are probabilities of "given that", ie:
p(b|a) = 1/2
p(a|b) = 4/5
→ Use Bayes theorem:
p(b)p(a|b) = p(a)p(b|a)
→ p(b) = (p(a)p(b|a))/p(a|b)
= (2/3 × 1/2) / (4/5)
= 2/3 × 1/2 × 5/4 = 5/12